OpenStudy (anonymous):

a = ?, b = 11, c = 17 A.) 12.96 B.) 12.75 C.) 13.20 D.) 13.12

OpenStudy (anonymous):

a^2 + b^2 = c^2 a^2 + 11^2 = 17^2 a^2 + 121 = 289 a^2 = 289 - 121 a^2 = 168 sq rt a^2 = sq rt 168 a = 12.96 just remember, if you get something like a^2 = 168...by taking the sq rt of both sides, it cancels out the ^2...leaving you with just a = sq rt 168

OpenStudy (anonymous):

and always make the hypotenuse (c) the biggest number